OpenStudy (anonymous):

Simplify 5/6(6x-12y)-1/4(8x-4y)

OpenStudy (anonymous):

\[=5(x-2y)-1(2x-y)\]\[=5x-10y-2x+y\]\[=3x-9y\]\[3(x-3y)\]

OpenStudy (anonymous):

Thanks!